OpenStudy (anonymous):

How many different basketball teams of 5 players can be created from a group of 7 players?

OpenStudy (anonymous):

35?

OpenStudy (anonymous):

no its C(7,5) = 7*6/2 = 21

OpenStudy (anonymous):

In how many ways can the letters of the word CARPET be arranged?

OpenStudy (anonymous):

I'm bad with these kinds of questions.

OpenStudy (anonymous):

6! = 720

OpenStudy (anonymous):

Last question, if you don't mind. A die with a 10% chance of rolling a 6 is rolled 10 times in a row. What is the probability that a 6 is rolled exactly 2 times?

OpenStudy (anonymous):

here n= 10 p= .1 hence q= 1-p = 1-0.1 =0.9 here we need f(2) (required 2 successes) C(10,2)*(0.1)^2*(0.9)^8 [C(10,2) = 45 ]

OpenStudy (anonymous):

45 isn't one of my options.

OpenStudy (anonymous):

only C(10,2) = 45